OpenStudy (alesia/bailey):

coffee: some say to brew an excellent cup of coffee, you must have a brewing temp. of 200 degrees F +/- 5 degrees write and solve an equation. max. and min. temps.

OpenStudy (triciaal):

let the temp = x ideal = 200 +/-5 then x = 200 + 5 or x = 200-5